OpenStudy (anonymous):

A student takes 40 units to 100 in a test. Teacher announces that this test and the next test wheels are averaged together for quality. The student wonders if caould get a C if it gets to 100 for the next test. He has reasons "I think my average (mean) is 70 because half of 40 is 20 and half of 100 is 50. This is a C because 20 plus 50 is 70. "Does the methood always work? explain your reasoning.

OpenStudy (anonymous):

see, i actually didn't get the question but still i am trying to solve it. apologize me if i may go wrong. he gets 40 marks in first test out of 100 and he expects to get 100 marks out of 100 in the next test so the mean is (40+100)/2 = 70 = 40/2 +100/2 = 70 so yes this will always work as far as there will be only two tests :P but in case of more than two test he needs to do like this 40/3 +100/3 ++++ (mark in third test)/3

OpenStudy (anonymous):

sorry, "in case of three tests" i meant!!